OpenStudy (anonymous):

Find the area of a regular hexagon with the given measurement. 6-inch radius. A = sq. in

OpenStudy (anonymous):

The sides of a regular hexagon are the same as the radius. - - - - - - - The area of any regular polygon with n sides of length s is: area=ns^2*cot(180/n)/4 = 6*36*cot(30)/4 = 54cot(30) = 54sqrt(3) =~ 93.53 sq in

OpenStudy (anonymous):

The answer looks like this tho ??sqrt??